# Personalized Pricing Through Strategic User Profiling in Social Networks

### v1

@@ I Introduction
Modern advances in information technology have enabled user profiling, which is a data analytic tool that outlines users’ characteristics like product preferences and demands. The availability of user profiling allows a seller to identify a user’s willingness to pay and exercise personalized pricing accordingly. For example, major online stores like Amazon and Sears track users’ web browsing history and online purchase records to personalize their offerings [2, 3]. Orbitz further differentiates users based on their computer’s operating systems and offers higher hotel prices to Mac users [4]. However, with the emergence of privacy-enhancing technologies (e.g., cookie blockers) that make user information no longer personally identifiable, today’s product sellers have encountered increasing obstacles when tracking users’ profiles only based on their on-site store data [5].
@@
Recent years have witnessed a rapid development of technologies that enable such user profiling with social network data [6]. For example, Facebook has launched APIs that enable several online stores’ access to users’ Facebook identity and profile information since 2008 [7]. Amazon and EyeBuyDirect have bought data from Facebook to train their personalization features [8, 9]. Apparently, such tracking and analysis of social network data can facilitate user profiling for product sellers to intensify price discrimination.
As various governments have recently tightened privacy regulations with a marked shift toward protecting users (e.g., [10, 11]), users are increasingly aware of such social data access and profiling practices from potential sellers. Particularly, the well-known informed consent policy requires online stores to inform people of their data access and the purpose of data usage. Following this, Amazon now needs to notify users of its social profile access in great detail when users sign in and connect to their Facebook accounts [9, 12]. Given such awareness, many users have reduced social media usage or even remain silent online to thwart the seller’s attempts to profile and exercise personalized pricing [13, 14]. We are thus motivated to ask the following two key questions.
@@
Addressing these questions presents several key challenges. Firstly, there exists a coupling between the seller and the users, as well as a coupling among the users themselves. These two types of couplings also interact with each other, which has not been systematically studied before. Regarding the first type of coupling, users are aware of the profiling risks when interacting with each other online and may adjust their interactions accordingly. Hence, the seller needs to carefully balance the uniform pricing for non-profiled users and the personalized pricing for profiled users to maximize profit. Previous works on traditional user profiling (e.g., [15, 16, 17]) viewed users as isolated individuals and analyzed the effect of each individual’s endeavor to conceal private information and bypass the seller’s profiling independently. However, these works ignored the coupling among users and, more specifically, the positive network externality of users’ information revelation. Thus, the insights gained from previous studies may not be applicable in this context, and this externality will significantly complicate the analysis of the implications of user profiling.
Another major challenge is the uncertainty involved in user profiling. Even if a user has exposed his personal information online, whether the seller can successfully profile this user is still probabilistic due to technological or regulatory constraints (e.g., [10, 11]). On the technical level, updating the seller’s beliefs about users’ private information based on user profiling results is subject to inherent uncertainty, as the profiling results may not directly relate to users’ online social behaviors. Therefore, this randomness in user profiling makes it more difficult to ensure consistency between users’ revealed information and the seller’s belief for further pricing. Note that satisfying such belief consistency is fundamental in analyzing interactions in a dynamic Bayesian setting [18].
@@
  • Tradeoff involving breaking a pricing promise: We further examine some prevalent pricing practices in our setting, where the seller breaks a pricing promise to personalize final offerings [19, 20]. Our analysis demonstrates that while this practice allows for the seller’s proactive control of users’ incentives in social decisions, it only marginally improves the seller’s average revenue while introducing higher variance. Considering the additional reputation and regulation risks [21], it is not recommended for sellers to break pricing promises in practice.
@@ II Related Work
A growing literature on personalized pricing has studied users’ purchase behaviors in online markets while considering the implications of revealing their private information to product sellers (e.g., [16, 17, 22, 23, 20, 24]). For example, Conitzer et al. in [16] investigated a repeated purchases scenario, where users could hide their past purchase records to hinder personalized pricing. Koh et al. in [22] further allowed users’ voluntary participation in profiling, while considering the benefit of reducing search costs for ideal products through users’ information revelation. Valletti et al. in [23] analyzed how users reveal or conceal their private information, given the seller’s strategic investment in consumer profiling technology. Chen et al. in [20] explored a duopoly setting, where sellers compete to target users and attempt to price discriminate while users make strategic purchase decisions to bypass.
@@
Several studies in the literature on personalized pricing have adopted PBE as the solution concept (e.g., [22, 23, 25, 16]). The standard analysis for PBE proceeds in three steps as follows. First, a belief about the structure of certain players’ strategies is proposed. Based on this proposed belief, the analysis for the remaining players’ strategies is then conducted through backward induction. Finally, the equilibrium strategies and the proposed belief are verified to constitute a PBE satisfying belief consistency and sequential rationality. However, the coupling among users’ social behaviors in our work significantly complicates the belief proposal in the first step. This departs from previous studies (e.g., [16, 17, 22]) that analyze users’ online behaviors individually, which allows for an easily identifiable threshold structure for the equilibrium belief. Motivated by the advanced game theory literature [26], we propose to explicitly conduct a forward analysis of users’ coupled social interactions to derive the structure of the equilibrium belief. In a nutshell, we alternate backward induction with forward induction to analyze the whole dynamic Bayesian game and provide a complete PBE characterization.
@@ III System Model
Consider the coupled interactions among the seller and a group of users $\mathcal{N}\triangleq\{1,\dots,n\}$ under user profiling. Specifically, users share information and interact with each other in the social network, while the shared social activity data is monitored or even accessible by the seller (e.g., through data-sharing deals with the social network platform [27]). Then, the seller manages to extract users’ private profiles from their online social activities to enable personalized pricing. [^1] For convenience, we hereafter use female pronouns to refer to the seller and male (or plural) pronouns to refer to each user (or users).
@@ III-A1 Social Interactions in Stage I
Each user $i$ gains more satisfaction in the social network if the social activity levels (of his and other users’) are higher. For instance, users feel connected, gain empathy and identity, and strengthen communication through more social interactions online [28]. Formally, we model the social network benefit $J_{i}$ for each user $i$ as below,
@@
which captures the network effect. Namely, the other users’ social activities impose a positive externality on one user’s social network benefit. The choice of the logarithmic function is motivated by the Zipf’s law [29], which emphasizes that the marginal network benefit each user $i$ experiences is diminishing as the others intensify their social activities.
@@ III-B Seller’s Model with User Profiling
The seller employs the profiling technology (e.g., [30, 31]) to analyze users’ social activity data and statistically infer their valuations (if possible). The user profiling signals the seller about one user’s valuation for the product or service. Based on the profiling results, the seller determines the prices (either personalized or uniform) offered to different users.
@@ III-B1 User Profiling
Similar to [17, 22, 23], we consider a typical binary profiler that outputs a signal about users’ valuation if and only if it is sure about the result. We assume that the profiling result about user $i$ follows a Bernoulli distribution. That is, user $i$’s valuation $v_{i}$ is revealed with a probability $\lambda_{i}\in[0,1]$; otherwise, user $i$ is non-profiled with unrevealed valuation. Here, the signal accuracy $\lambda_{i}$ is increasing in user $i$’s social activity level $x_{i}$, i.e.,
@@
The signal accuracy in (6) is a concave function of $x_{i}$ with $0<\alpha<1$. It indicates that, as the seller already harvests lots of critical information, the extra data extracted from user $i$’s online social activity becomes less informative. Note that (6) is upper bounded by $\delta$, referred to as user profiling accuracy. This practically captures the fact that, due to technological and regulatory constraints (e.g., [10, 11]), whether the seller can successfully profile a user is probabilistic even when this user is fully active in the social network.
One interpretation of such profiler assumption is to consider a binary hypothesis testing of user $i$’s valuation at value $v_{i}$ with a confidence level $\lambda_{i}$ [23]. Our analysis method also applies to more general profiling models. For instance, we can extend our analysis to the case with noisy profiling signals, where a user $i$’s valuation $v_{i}$ is revealed as a random variable over the support $a_{i},b_{i}]$ following a cumulative distribution function $G(\cdot)$. Instead of choosing the price $p_{i}^{*}=v_{i}$ in (1), the seller in this more general case can sets the optimal personalized price as $p_{i}^{*}=\arg\max_{p_{i}\in[a_{i},b_{i}]}p_{i}(1-G(p_{i}))$. For clarity, we focus on the binary profiler to crystalize insights in this work.
@@ III-C Dynamic Bayesian Game Formulation
Due to the double-coupled decision-making of users’ social activities and the seller’s (uniform) pricing in Stages I and II, backward induction alone cannot solve the equilibrium. To ensure belief consistency over stages, we further combine forward induction motivated by the advanced game theory literature [26]. That is, we alternate backward induction with forward induction to analyze the PBE of our dynamic Bayesian game (see Section IV for details).
@@ IV-B Forward Analysis of Users’ Social Decisions in Stage I
  • Step I (Game Formulation and Solution Existence). We first identify users’ social interactions in Stage I as a static Bayesian game with strategic complementarities [32], where users make their social activity decisions while predicting the seller’s uniform price $p_{0}$. This then indicates the existence of a pure-strategy Bayesian Nash equilibrium in users’ social activity levels.
@@ V Analysis of PBE
After alternating backward analysis with forward analysis in Stages I and II, we now derive the PBE of the whole dynamic Bayesian game in this section. Specifically, we combine users’ equilibrium valuation threshold $v^{\ast}$ for their social activity decisions (see Theorem 1) and the seller’s optimal uniform pricing scheme $p_{0}^{*}$ (see Proposition 1) together. In this way, we meet the two requirements of PBE: (i) sequential rationality: the seller sets the uniform price to maximize her sale revenue in Stage II given her belief of users’ social activities, which is derived through Bayes’ theorem; (ii) belief consistency: users make their social activity decisions in Stage I while predicting the seller’s pricing in Stage II, and such equilibrium strategies should be consistent with the seller’s belief [33].
@@ VI-A Effect on Users’ Information Disclosure
In the no-social-network benchmark, there exists a unique user-optimal PBE as follows. [^6] Notice that multiple equilibria exist in this no-social-network benchmark, and we will focus on the user-optimal equilibrium to provide clear insights, which maximizes the user’s ex-ante payoff across all equilibria. This refinement, known as the sender-optimal equilibrium, is widely adopted in the literature on incomplete information games (e.g., [34]).
@@ VI-B Performance Evaluation for the Seller
Next, we evaluate the performance of our proposed pricing mechanism for the seller, through which we will understand the power of harnessing social network benefits. To evaluate the performance, we will introduce two additional benchmarks adapted from the existing literature (e.g., [16, 22, 35, 25]). Due to space constraints, we have deferred the detailed descriptions of the pricing benchmarks and the performance evaluation to Appendix O.
@@ VII Impact of Pricing Timing
Thus far, we have assumed that the seller determines the uniform and personalized prices simultaneously in Stage II of the three-stage model in Section III. However, some sellers currently employ alternative pricing practices where they initially make a pricing promise but break it later to personalize final offerings [19, 20]. This raises the issue of whether the seller should adopt such promise-breaking pricing practices while profiling users in the social network. To address this concern, in this section, we further consider the seller’s flexibility in the pricing timing, i.e., the seller can announce the uniform price before users’ social activity decisions. Such a pricing practice will make the uniform price publicly observable to all users. Notice that personalized pricing is always determined after user profiling and offered to the profiled users privately, the same as our original model.
@@
As the seller provides personalized prices after announcing the uniform price, she actually breaks the pricing commitment to some users. This may harm the seller’s reputation or even violate related regulations on pricing fraud [21]. Later in this section, we will show the surprising result: Even without considering reputation or regulation issues, such flexibility in pricing timing only marginally improves the seller’s average revenue but at the cost of higher variance. Therefore, it is not reasonable for the seller to adopt a flexible pricing timing approach. This further justifies our use of the original model in Section III as the focus of this paper.
@@ VII-B Impact of Flexibility in Pricing Timing on Seller’s Revenue
Finally, we turn our attention to the variance in the seller’s revenue. Notice that examining revenue variance could assist the seller’s risk assessment [36] by telling the potential revenue volatility that mainly stems from randomness in user profiling. We conduct extensive simulation experiments to compare the actual revenue distributions in the two models. As shown in Fig. 10, the black dashed and the red solid curves represent the seller’s average revenues of our original three-stage model and the extended four-stage model, respectively. We also plot vertical error bars across the curves to illustrate the revenue variances. Moreover, we graphically depict the improvement ratio of the seller’s average revenue and the increment ratio of the seller’s revenue variance in the extended model compared to our original model through blue and green curves in Fig. 10, respectively. Actually, we note that the seller can only improve the revenue by no more than $2\%$ with her flexibility in pricing timing.

### v3

@@ Abstract
Index Terms:
Online social networks, user profiling technology, privacy, personalized pricing, dynamic Bayesian game.
@@ I Introduction
Modern advances in information technology have enabled user profiling, which is a data analytic tool that outlines users’ characteristics like product preferences and demands. The availability of user profiling allows a seller to identify a user’s willingness to pay and exercise personalized pricing accordingly. For example, major online stores like Amazon and Sears track users’ web browsing history and online purchase records to personalize their offerings [howe_a, datadriven]. Orbitz further differentiates users based on their computer’s operating systems and offers higher hotel prices to Mac users [mattioli_2012]. However, with the emergence of privacy-enhancing technologies (e.g., cookie blockers) that make user information no longer personally identifiable, today’s product sellers have encountered increasing obstacles when tracking users’ profiles only based on their on-site store data [anant_2020_consumer].
@@
Recent years have witnessed a rapid development of technologies that enable such user profiling with social network data [2019Social]. For example, Facebook has launched APIs that enable several online stores’ access to users’ Facebook identity and profile information since 2008 [announcing]. Amazon and EyeBuyDirect have bought data from Facebook to train their personalization features [bustos_2011_facebook, bustos_2011_7]. Apparently, such tracking and analysis of social network data can facilitate user profiling for product sellers to intensify price discrimination.
As various governments have recently tightened privacy regulations with a marked shift toward protecting users (e.g., [a2013_general, a2018_california]), users are increasingly aware of such social data access and profiling practices from potential sellers. Particularly, the well-known informed consent policy requires online stores to inform people of their data access and the purpose of data usage. Following this, Amazon now needs to notify users of its social profile access in great detail when users sign in and connect to their Facebook accounts [bustos_2011_7, boulton_2010_amazon]. Given such awareness, many users have reduced social media usage or even remain silent online to thwart the seller’s attempts to profile and exercise personalized pricing [cisco_2019_cisco, daxtheduck_2019_new]. We are thus motivated to ask the following two key questions.
@@
Addressing these questions presents several key challenges. Firstly, there exists a coupling between the seller and the users, as well as a coupling among the users themselves. These two types of couplings also interact with each other, which has not been systematically studied before. Regarding the first type of coupling, users are aware of the profiling risks when interacting with each other online and may adjust their interactions accordingly. Hence, the seller needs to carefully balance the uniform pricing for non-profiled users and the personalized pricing for profiled users to maximize profit. Previous works on traditional user profiling (e.g., [acquisti2005conditioning, conitzer2012hide, 2015Monopoly]) viewed users as isolated individuals and analyzed the effect of each individual’s endeavor to conceal private information and bypass the seller’s profiling independently. However, these works ignored the coupling among users and, more specifically, the positive network externality of users’ information revelation. Thus, the insights gained from previous studies may not be applicable in this context, and this externality will significantly complicate the analysis of the implications of user profiling.
Another major challenge is the uncertainty involved in user profiling. Even if a user has exposed his personal information online, whether the seller can successfully profile this user is still probabilistic due to technological or regulatory constraints (e.g., [a2013_general, a2018_california]). On the technical level, updating the seller’s beliefs about users’ private information based on user profiling results is subject to inherent uncertainty, as the profiling results may not directly relate to users’ online social behaviors. Therefore, this randomness in user profiling makes it more difficult to ensure consistency between users’ revealed information and the seller’s belief for further pricing. Note that satisfying such belief consistency is fundamental in analyzing interactions in a dynamic Bayesian setting [fudenberg1991perfect].
@@
  • Tradeoff involving breaking a pricing promise: We further examine some prevalent pricing practices in our setting, where the seller breaks a pricing promise to personalize final offerings [thecouncilofeconomicadvisers_2015_the, chen2020competitive]. Our analysis demonstrates that while this practice allows for the seller’s proactive control of users’ incentives in social decisions, it only marginally improves the seller’s average revenue while introducing higher variance. Considering the additional reputation and regulation risks [lee_2021_chinese], it is not recommended for sellers to break pricing promises in practice.
@@ II Related Work
A growing literature on personalized pricing has studied users’ purchase behaviors in online markets while considering the implications of revealing their private information to product sellers (e.g., [conitzer2012hide, 2015Monopoly, 2017Is, 2020Consumer, chen2020competitive, bimpikis2021data]). For example, Conitzer et al. in [conitzer2012hide] investigated a repeated purchases scenario, where users could hide their past purchase records to hinder personalized pricing. Koh et al. in [2017Is] further allowed users’ voluntary participation in profiling, while considering the benefit of reducing search costs for ideal products through users’ information revelation. Valletti et al. in [2020Consumer] analyzed how users reveal or conceal their private information, given the seller’s strategic investment in consumer profiling technology. Chen et al. in [chen2020competitive] explored a duopoly setting, where sellers compete to target users and attempt to price discriminate while users make strategic purchase decisions to bypass.
@@
Several studies in the literature on personalized pricing have adopted PBE as the solution concept (e.g., [2017Is, 2020Consumer, 2020Voluntary, conitzer2012hide]). The standard analysis for PBE proceeds in three steps as follows. First, a belief about the structure of certain players’ strategies is proposed. Based on this proposed belief, the analysis for the remaining players’ strategies is then conducted through backward induction. Finally, the equilibrium strategies and the proposed belief are verified to constitute a PBE satisfying belief consistency and sequential rationality. However, the coupling among users’ social behaviors in our work significantly complicates the belief proposal in the first step. This departs from previous studies (e.g., [conitzer2012hide, 2015Monopoly, 2017Is]) that analyze users’ online behaviors individually, which allows for an easily identifiable threshold structure for the equilibrium belief. Motivated by the advanced game theory literature [battigalli2006rationalization], we propose to explicitly conduct a forward analysis of users’ coupled social interactions to derive the structure of the equilibrium belief. In a nutshell, we alternate backward induction with forward induction to analyze the whole dynamic Bayesian game and provide a complete PBE characterization.
@@ III System Model
Consider the coupled interactions among the seller and a group of users $\mathcal{N}\triangleq\{1,\dots,n\}$ under user profiling. Specifically, users share information and interact with each other in the social network, while the shared social activity data is monitored or even accessible by the seller (e.g., through data-sharing deals with the social network platform [x_2018]). Then, the seller manages to extract users’ private profiles from their online social activities to enable personalized pricing. [^1] For convenience, we hereafter use female pronouns to refer to the seller and male (or plural) pronouns to refer to each user (or users).
@@ III-A1 Social Interactions in Stage I
Each user $i$ gains more satisfaction in the social network if the social activity levels (of his and other users’) are higher. For instance, users feel connected, gain empathy and identity, and strengthen communication through more social interactions online [seiter2015secret]. Formally, we model the social network benefit $J_{i}$ for each user $i$ as below,
@@
which captures the network effect. Namely, the other users’ social activities impose a positive externality on one user’s social network benefit. The choice of the logarithmic function is motivated by the Zipf’s law [2006Metcalfe], which emphasizes that the marginal network benefit each user $i$ experiences is diminishing as the others intensify their social activities.
@@ III-B Seller’s Model with User Profiling
The seller employs the profiling technology (e.g., [2013Private, 2018Collective]) to analyze users’ social activity data and statistically infer their valuations (if possible). The user profiling signals the seller about one user’s valuation for the product or service. Based on the profiling results, the seller determines the prices (either personalized or uniform) offered to different users.
@@ III-B1 User Profiling
Similar to [2015Monopoly, 2017Is, 2020Consumer], we consider a typical binary profiler that outputs a signal about users’ valuation if and only if it is sure about the result. We assume that the profiling result about user $i$ follows a Bernoulli distribution. That is, user $i$’s valuation $v_{i}$ is revealed with a probability $\lambda_{i}\in[0,1]$; otherwise, user $i$ is non-profiled with unrevealed valuation. Here, the signal accuracy $\lambda_{i}$ is increasing in user $i$’s social activity level $x_{i}$, i.e.,
@@
The signal accuracy in (6) is a concave function of $x_{i}$ with $0<\alpha<1$. It indicates that, as the seller already harvests lots of critical information, the extra data extracted from user $i$’s online social activity becomes less informative. Note that (6) is upper bounded by $\delta$, referred to as user profiling accuracy. This practically captures the fact that, due to technological and regulatory constraints (e.g., [a2013_general, a2018_california]), whether the seller can successfully profile a user is probabilistic even when this user is fully active in the social network.
One interpretation of such profiler assumption is to consider a binary hypothesis testing of user $i$’s valuation at value $v_{i}$ with a confidence level $\lambda_{i}$ [2020Consumer]. Our analysis method also applies to more general profiling models. For instance, we can extend our analysis to the case with noisy profiling signals, where a user $i$’s valuation $v_{i}$ is revealed as a random variable over the support $a_{i},b_{i}]$ following a cumulative distribution function $G(\cdot)$. Instead of choosing the price $p_{i}^{*}=v_{i}$ in (1), the seller in this more general case can sets the optimal personalized price as $p_{i}^{*}=\arg\max_{p_{i}\in[a_{i},b_{i}]}p_{i}(1-G(p_{i}))$. For clarity, we focus on the binary profiler to crystalize insights in this work.
@@ III-C Dynamic Bayesian Game Formulation
Due to the double-coupled decision-making of users’ social activities and the seller’s (uniform) pricing in Stages I and II, backward induction alone cannot solve the equilibrium. To ensure belief consistency over stages, we further combine forward induction motivated by the advanced game theory literature [battigalli2006rationalization]. That is, we alternate backward induction with forward induction to analyze the PBE of our dynamic Bayesian game (see Section IV for details).
@@ IV-B Forward Analysis of Users’ Social Decisions in Stage I
  • Step I (Game Formulation and Solution Existence).We first identify users’ social interactions in Stage I as a static Bayesian game with strategic complementarities [1990Rationalizability], where users make their social activity decisions while predicting the seller’s uniform price $p_{0}$. This then indicates the existence of a pure-strategy Bayesian Nash equilibrium in users’ social activity levels.
@@ V Analysis of PBE
After alternating backward analysis with forward analysis in Stages I and II, we now derive the PBE of the whole dynamic Bayesian game in this section. Specifically, we combine users’ equilibrium valuation threshold $v^{\ast}$ for their social activity decisions (see Theorem 1) and the seller’s optimal uniform pricing scheme $p_{0}^{*}$ (see Proposition 1) together. In this way, we meet the two requirements of PBE: (i) sequential rationality: the seller sets the uniform price to maximize her sale revenue in Stage II given her belief of users’ social activities, which is derived through Bayes’ theorem; (ii) belief consistency: users make their social activity decisions in Stage I while predicting the seller’s pricing in Stage II, and such equilibrium strategies should be consistent with the seller’s belief [fudenberg1991game].
@@ VI-A Effect on Users’ Information Disclosure
In the no-social-network benchmark, there exists a unique user-optimal PBE as follows. [^6] Notice that multiple equilibria exist in this no-social-network benchmark, and we will focus on the user-optimal equilibrium to provide clear insights, which maximizes the user’s ex-ante payoff across all equilibria. This refinement, known as the sender-optimal equilibrium, is widely adopted in the literature on incomplete information games (e.g., [mailath1993belief]).
@@ VI-B Performance Evaluation for the Seller
Next, we evaluate the performance of our proposed pricing mechanism for the seller, through which we will understand the power of harnessing social network benefits. To evaluate the performance, we will introduce two additional benchmarks adapted from the existing literature (e.g., [conitzer2012hide, 2017Is, zhang2011perils, 2020Voluntary]). Due to space constraints, we have deferred the detailed descriptions of the pricing benchmarks and the performance evaluation to Appendix O.
@@ VII Impact of Pricing Timing
Thus far, we have assumed that the seller determines the uniform and personalized prices simultaneously in Stage II of the three-stage model in Section III. However, some sellers currently employ alternative pricing practices where they initially make a pricing promise but break it later to personalize final offerings [thecouncilofeconomicadvisers_2015_the, chen2020competitive]. This raises the issue of whether the seller should adopt such promise-breaking pricing practices while profiling users in the social network. To address this concern, in this section, we further consider the seller’s flexibility in the pricing timing, i.e., the seller can announce the uniform price before users’ social activity decisions. Such a pricing practice will make the uniform price publicly observable to all users. Notice that personalized pricing is always determined after user profiling and offered to the profiled users privately, the same as our original model.
@@
As the seller provides personalized prices after announcing the uniform price, she actually breaks the pricing commitment to some users. This may harm the seller’s reputation or even violate related regulations on pricing fraud [lee_2021_chinese]. Later in this section, we will show the surprising result: Even without considering reputation or regulation issues, such flexibility in pricing timing only marginally improves the seller’s average revenue but at the cost of higher variance. Therefore, it is not reasonable for the seller to adopt a flexible pricing timing approach. This further justifies our use of the original model in Section III as the focus of this paper.
@@ VII-B Impact of Flexibility in Pricing Timing on Seller’s Revenue
Finally, we turn our attention to the variance in the seller’s revenue. Notice that examining revenue variance could assist the seller’s risk assessment [baron2022revenue] by telling the potential revenue volatility that mainly stems from randomness in user profiling. We conduct extensive simulation experiments to compare the actual revenue distributions in the two models. As shown in Fig. 10, the black dashed and the red solid curves represent the seller’s average revenues of our original three-stage model and the extended four-stage model, respectively. We also plot vertical error bars across the curves to illustrate the revenue variances. Moreover, we graphically depict the improvement ratio of the seller’s average revenue and the increment ratio of the seller’s revenue variance in the extended model compared to our original model through blue and green curves in Fig. 10, respectively. Actually, we note that the seller can only improve the revenue by no more than $2\%$ with her flexibility in pricing timing.